OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After three hours, the velocity of the car is 51 km/h. After five hours, the velocity of the car is 59 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. Part B: How can you graph the equation obtained in Part A for the first six hours?

OpenStudy (anonymous):

3 hours--------51m/s 5hours ---------59m/s ignore the word, just focus on the numbers, you have 2 points (3, 51) and (5,59) . to write an equation of a line, you need : 1/ slope. (can get it from 2 points) 2/ 1 point . 1/ SLOPE: We have a formula to find it from 2 given points, it says: \[m = \frac{ y _{2} - y _{1} }{ x _{2 - x _{1}} }\] can you show me how to apply our 2 points to this formula?

OpenStudy (anonymous):

would you \[M= 3,51\over 5,59\]

OpenStudy (anonymous):

not m up top tho

OpenStudy (anonymous):

59 - 51 ------- 5 - 3 8 -- 2 4 -- 1 The slope is 4

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

so the slope is 4 and the y intercept is 51/59

OpenStudy (anonymous):

ok, the last step is put slope, point into formula

OpenStudy (anonymous):

y=4+51/59 ?

OpenStudy (anonymous):

\[y - 51 = 4(x-3)\]

OpenStudy (anonymous):

simplify them, isolate y, what do you get?

OpenStudy (anonymous):

y=4x + 63

OpenStudy (anonymous):

redo, -12 +51≠63

OpenStudy (anonymous):

39

OpenStudy (anonymous):

it's y = 4x + 39, now, back to your problem. we should choose v stands for velocity instead of y, and t for hour instead of x, right? so, our equation should be v = 4t +39

OpenStudy (anonymous):

yay so part a is over right :D

OpenStudy (anonymous):

Yes, that is it for Part A. Next you just sketch the graph, do you know how to sketch the graph?

OpenStudy (anonymous):

You don't graph it tho you say how you would i thought

OpenStudy (anonymous):

Oh sorry, I meant to say that :3

OpenStudy (anonymous):

:D so i would say sub 6 for t

OpenStudy (anonymous):

Plot the y intercept, then use the slope. the easiest way is: let x =0, find y let y =0, find x link 2 points (0,y) and (x,0) to get the graph.

OpenStudy (anonymous):

Now you are all done! Thank you for letting me help you, Signed: \[Mangorox\]

OpenStudy (anonymous):

THANSKS<3